OpenStudy (anonymous):

An arithmetic sequence is represented by the explicit formula A(n) = 2 + 9(n - 1). What is the recursive formula? A. A(n) = A(n - 1) + 2 B. A(n) = A(n - 1) + 9 C. A(n) = A(n - 1) - 9 D. A(n) = A(n - 1) - 2

hartnn (hartnn):

A(n-1) = 2+ (n-2)*9 = [2+ (n-1)*9] - 9 = ??

OpenStudy (anonymous):

I don't know.

hartnn (hartnn):

A(n-1) = 2+ (n-2)*9 = [2+ (n-1)*9] - 9 = A(n)-9 so A(n) = A(n-1) + 9.

OpenStudy (anonymous):

Ah, I see.

hartnn (hartnn):

did u understand the adjustments i did?

OpenStudy (anonymous):

Yes, thank you.

hartnn (hartnn):

welcome :)